OpenStudy (anonymous):

just two questions i need help with Which of the following ordered pairs is not a solution to the inequality y −4x − 2?

OpenStudy (anonymous):

(−1, 0) (0, 0) (4, 4) (1, 3)

OpenStudy (anonymous):

Which of the following ordered pairs is a solution to the inequality y ≥4x − 2? (0, −4) (0, 0) (0, −3) (0, −5)

OpenStudy (anonymous):

first one is actually Which of the following ordered pairs is not a solution to the inequality y ≥ −4x − 2?

OpenStudy (anonymous):

please help anyone!

OpenStudy (anonymous):

Check each ordered pair by substituting the values of x and y into the inequality. I'll check the first one for you. y >= -4x - 2 (−1, 0) 0 >= -4(-1) - 2 0 >= 4 - 2 0 >= 2 Not true^ Can you do that for the rest?

OpenStudy (anonymous):

no not really because i have no clue about inequalities

OpenStudy (anonymous):

like the site doesnt explain it

OpenStudy (anonymous):

wait is it -1,0

OpenStudy (anonymous):

y >= -4x - 2 is an inequality. Do you understand what I did here? Substitute the ordered pairs (x, y) into the inequality. (x, y) (−1, 0) y >= -4x - 2 0 >= -4(-1) - 2 0 >= 4 - 2 0 >= 2 0 is NOT greater than or equal to(>=) 2. SO (-1, 0) is wrong.

OpenStudy (anonymous):

and is the 2nd 0,-4

OpenStudy (anonymous):

y ≥ 4x − 2 -4 ≥ 4(0) - 2 -4 ≥ 0 - 2 -4 ≥ -2 -4 is NOT greater than/ equal to -2. So it's incorrect.

OpenStudy (anonymous):

@ilovejesus316 I lve you name @ilovejesus316